OpenStudy (sloppycanada):

What is the slope of the line that is perpendicular to y = 2x – 3? How do I do this? Please don't give me the answer.

OpenStudy (anonymous):

y=mx+b For two lines to be perpendicular, their two slopes (m) must be -1 when it gets multiplied. y=2x-3 2 represemts m in y=mx+b, so the slope is 2 2*(slope of other line)=-1

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Y=(slope)x+y-intercept thats why

OpenStudy (sloppycanada):

Okay so, y = 3x – 3 the slope would be 1/3?

OpenStudy (anonymous):

no its just 3

OpenStudy (anonymous):

or 3/1

OpenStudy (anonymous):

Slope of which line? y=3x-3 or the line perpendicular to y=3x-3? Btw dont listen to the person that posted after me. It is not correct.

jimthompson5910 (jim_thompson5910):

take the original slope flip the fraction and flip the sign to get the perpendicular slope

jimthompson5910 (jim_thompson5910):

so for example, if we had the original slope to be 2/7, then the perpendicular slope would be -7/2 another example original: -4/9 perpendicular: 9/4 these are just examples and none of this is the answer

OpenStudy (sloppycanada):

wait. it wants the parallel to y = 3x - 3

jimthompson5910 (jim_thompson5910):

parallel lines have equal slopes

OpenStudy (anonymous):

How are you skipping to another equation and going from perpendicular to parallel? Are you done with your first question?

OpenStudy (sloppycanada):

okay so it'd be three

OpenStudy (sloppycanada):

?

OpenStudy (anonymous):

yup

OpenStudy (sloppycanada):

the slope of the first one, the perpendicular thing, would be -2 since apparently it's the opposite of the original.

jimthompson5910 (jim_thompson5910):

sloppycanada, you forgot to flip the fraction

jimthompson5910 (jim_thompson5910):

think of 2 as 2/1

OpenStudy (sloppycanada):

So it'd be -1/2? Since it's positive in the original?

jimthompson5910 (jim_thompson5910):

yeah if +2/1 is the original slope, then -1/2 is the perpendicular slope

OpenStudy (sloppycanada):

What is the slope of the line that is parallel to y = 3x – 3? so for this it'd just be 3?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (sloppycanada):

Thanks a bunch!